OpenStudy (anonymous):

college algebra, CANNOT figure this out: a 2 inch by 4 inch by 7 inch iron block is plated with brass so that each of its three dimensions is increased by the same amount x. the volume of the plated block is 50% greater than the volume of the unplated block. then it asks to give the function for the volume of the plated block. and i know that is (2+x)(4+x)(7+x) and then it asks what are the three dimensions of the plated block. and the answers are 2.4941, 4.4941, and 7.4941.

OpenStudy (mathmate):

So you need to solve for the thickness x, such that 1.5(2*4*7) = (2+x)(4+x)(7+x) or f(x)=(2+x)(4+x)(7+x)-84=0 Graphing tells us that there is a root around 0.5. Use Newton's method to converge to the root: f'(x)=3x^2+26x+50, x=0.5-f(0.5)/f'(0.5)=0.4941

OpenStudy (anonymous):

thank you so so much! like you have no idea. thank you!

OpenStudy (mathmate):

You're welcome!